OpenStudy (anonymous):

Evaluate: Sigma(Infinite) i = 1 Formula: 3(-1/4)^i (sum)

OpenStudy (anonymous):

Okay maybe this \[\sum_{i=1}^{\infty} 3 \times \frac{-1}{4}^i \]

OpenStudy (anonymous):

precisely

OpenStudy (anonymous):

it just says: Evaluate next to it

OpenStudy (anonymous):

okay do you know sum of gp (geometric progression )

OpenStudy (anonymous):

yes. s = 1/(1-r)

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

thats correct :)

OpenStudy (anonymous):

\[Sum_{G.P} = \frac{a}{1-r}\] \[r = \frac{-1}{4}\] \[a = \frac{-1}{4}\] \[Sum = 3 ( \frac{ \frac{-1}{4}}{1 + \frac{1}{4} })\]

OpenStudy (anonymous):

thank u. i suppose i was right and entered things wrong.

OpenStudy (anonymous):

\[Sum = -\frac{3}{5}\]

OpenStudy (anonymous):

: )